Greenhab Report – April 2nd

GreenHab Report
Kai Takeda
2 APR 2018

Environmental control: (Choose which of the following is appropriate and explain further if needed)

Ambient with window/door open
Shade cloth on
Fan on(10:00~15:00)

Average temperatures: 24.0℃
Low 14.6℃
High 40.3℃

Hours of supplemental light: None

Changes to crops: None

Daily water usage for crops: 4gal

Time(s) of watering for crops: twice

Moringa research observations: None

Changes to research plants: No

Aquaponics: None

Narrative:
Dear Dr.Shannon
Thank you for your e-mail. I follow your instruction.
Should I thin out seedlings which I sow last week (radish sprouts and lettuce and so on) ?

Support/supplies needed: None

Operations Report – April 2nd

Crew 191 Operations Report 02/04/2018

SOL: 9

Name of person filing report: Wataru Okamoto

Non-nominal systems: Solar system, Water Tank
Notes on non-nominal systems: Solar system seems still unusual from Sol5. But Solar system condition better than yesterday.Please reference under General notes. We collected sub tank contamination.

Generator (hours run): 12hours

Solar SOC 91% (Before generator is run at night)

Diesel Almost FULL

Propane 72psi.

Ethanol Free Gasoline (5 Gallon containers for ATV) ? 6gallons.

Water (sub) –

Water (static) 90%

Trailer to Static Pump used ? –

Water (loft) Static to Loft Pump used Yes

Water Meter: 132936.4
Toilet tank emptied: Yes

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3) –

Oil Added –

ATV Fuel Used: –
Hours the ATVs were Used today: –

Notes on ATVs: –

Deimos rover used: –
Hours: 117.0h
Beginning charge: 100%
Ending charge: 100%
Currently charging: Yes

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours:
Beginning charge:
Ending charge:
Currently charging:

Spirit rover used: –
Hours: 29.2h
Beginning charge: 100%
Ending charge: 100%
Currently charging: –

Opportunity rover used: –
Hours: 28.4h
Beginning charge: 100%
Ending charge: 100%
Currently charging: –

Curiosity rover used: –
Hours: 23.8h
Beginning charge: 100%
Ending charge: 100%
Currently charging: –

HabCar used and why, where? –

General notes and comments: Solar Battery became 100% at 9:13.We are still process of cleaning water contamination.

Summary of internet: Under control

Summary of suits and radios:

– We are checking radios sensitivity.
– We repaired No.2 Backpack Hose
– No.4 Backpack charging socket lock notch was broken.

Summary of Hab operations: No problem

Summary of GreenHab operations: No problem

Summary of ScienceDome operations:

Summary of RAMM operations: The sub water tank was carried to RAMM.

Summary of health and safety issues: No problem

Questions, concerns and requests to Mission Support:
No.4 Backpack charging socket lock notch was broken.
Please supply us the charging socket plate indicate attached photo. Thank you for Dr. Shannon support. We need another 5 gallons drinking water tomorrow.

EVA Report – April 2nd

EVA Report

EVA #10 – 2 Apr 2018 – Sol9
Crew members: Yusuke Murakami (EVA leader), Fumiei Morisawa, Makoto Kawamura

Objectives:
– Operational check
– Sample the water of the white tank and the black tank
– Empty the water of the white tank and move to the RAM

Operations:
10:37 – Entering the airlock
10:40 – Operational EVA start
11:17 – Entering into RAM’s Airlock
11:20 – Entering RAM
11:21 – Open the shutter of RAM
11:29 – Put the white tank inside RAM
11:36 – Entering the airlock of the HAB
11:39 – Entering the HAB

Narrative:
This morning we decided to cancel EVA#9 (Venzha’s CEMs Project) because we had the water tank problems. After the daily operational check, we checked and sampled the water inside the white tank and the black tank. Shannon had already pumped out almost all of the water this morning, we fully emptied the water from the white tank. After that we measured the size of the tank and also the shutter of the RAM. It seemed that the tank could be pass into the RAM through the shutter. So we tried it, and did it. So now we got a choice to clean the tank without breaking the SIM later on.

Vehicles
– none

Crew Photos – April 2nd

SOL Essay (Photo Description)
02 Apr 2018

Sol 09
Miho Tsukishiro (XO)

This morning we have decided our EVA #9, because we have to solve the water pollution in the HAB. Since we found some white unknown material in the white tank yesterday, we can not use the water anymore, which already in the HAB water system.
The solution we took without breaking our SIM: request supply of clean water for two days, and clean up the tanks meanwhile.
So we checked and got some sample of water from the tanks and brought the empty but still contaminated white tank into the RAM today, so that we can clean it up during following EVA.

Sol Summary – April 2nd

Crew 191 Sol Summary Report
02 Apr 2018

Sol 09

Summary Title:
Water trouble
Miho Tsukishiro (XO)

Mission Status:
in progress

Sol Activity Summary:
8:00 Briefing
10:30 EVA #10 for operational check and water trouble management (leader: Yusuke Murakami, Fumiei Morisawa, Makoto Kawamura) Habcom: Miho Tsukishiro
13:00 Supply of Water
13:30 Lunch
18:00 Gathering reports
19:00 CapCom

Look Ahead Plan:
RVA #9: CEM’s EVA (Venzha’s Project)

Anomalies in work:
N/A

Weather:
Partly cloudy. Strong winds.

Crew Physical Status:
Good.

EVA:
See EVA Report.

Reports to be file:
– SOL Summary Report
– Engineer Report
– GreenHab Report
– Photo Report with Description
– EVA Report
– EVA Request
– HSO Report

Support Requested:
N/A